You want to purchase Pepsi 12 packs for Thanksgiving Dinner. The average price of the 12...

  1. You want to purchase Pepsi 12 packs for Thanksgiving Dinner. The average price of the 12 pack is $6.98 with a standard deviation of $1.03. Calculate the appropriate sample size where you want to be 90% confident in your results and be within a margin of error of $.25 on the price of 12 packs.
Answer #1

Solution :

Given that,

standard deviation = \sigma =1.03

margin of error = E = 0.25

At 90% confidence level the z is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.10

\alpha / 2 = 0.10 / 2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

Sample size = n = ((Z\alpha/2 * \sigma ) / E)2

= ((1.645 * 1.03) / 50.25)2

= 46

Sample size = 46